3.9GHz status
Monday, 18 August 2008 -
09:00
Monday, 18 August 2008
09:00
Status of cavity 3 engineering note
Status of cavity 3 engineering note
09:00 - 09:30
Room: The Hermitage
09:30
A0 update -- HPR system -- vacuum system
A0 update -- HPR system -- vacuum system
09:30 - 09:50
Room: The Hermitage
09:50
Status of attempted repair of cavity 2
Status of attempted repair of cavity 2
09:50 - 10:05
Room: The Hermitage
10:05
Other business
Other business
10:05 - 10:20
Room: The Hermitage